OpenStudy (anonymous):

Solve the linear system of equations 9x-5y=2 y=2x+1

OpenStudy (anonymous):

try to put 2x to left side

OpenStudy (pinksapphire):

x=-7 y=-13

OpenStudy (anonymous):

how do you work out the problem.. PinkSapphire, thats what I need help with.

OpenStudy (pinksapphire):

Okay...

OpenStudy (pinksapphire):

9x-5y=2 Take one of the problems. 9x-5(2x+1)=2 Put 2x+1 in for y because of the other equation. 9x-10x-5=2 Distribute. -1x-5=2 Add 5 to both sides +5 +5 -1x/-1=7/-1 Divide by -1 to each side to get x by itself. x=-7 Now that you have x, plug it in for 9x-5y=2 9(-7)-5y=2 -63-5y=2 Add 63 to both sides. +63 +63 -5y=65 -5y/-5=65/-5 Divide by -5 on each side to get y by itself. y=-13

OpenStudy (anonymous):

place each equation in the same form, then multiply one or both equations so you are able to cancel out a variable. the remaining number is the answer for the remaining variable. go back to the original system of equations and substitute the number for the variable. the answer is then shown.

OpenStudy (pinksapphire):

Do you understand it?

OpenStudy (anonymous):

Yes thank you very much, you're a life saver.

OpenStudy (pinksapphire):

Okay :) Glad I can help :)